# Lattice Boltzmann for linear elastodynamics: periodic problems and Dirichlet boundary conditions

### v1

@@ 1 Introduction
To circumvent the aforementioned issues, in this paper we adopt the so-called vectorial lattice Boltzmann method, which involves vector-valued instead of scalar populations [24, 25, 26, 27, 28, 29]. In a sense, multi distribution formulations such as those in [30, 31] can be also subsumed into the class of vectorial LBM. If we consider the so-called acoustic scaling, it is straightforward to show that the vectorial LBM with a vectorial counterpart of the well-known BGK collision operator approximates first-order hyperbolic systems of PDEs with second-order consistency in the non-dissipative limit [25]. In fact, the vectorial LBM can be seen as a particularly simple numerical approximation of so-called relaxation schemes for systems of conservation laws [32, 33].
@@ 2 Linear elastodynamics and its reformulation for LBM
In this section, we start by introducing our target problem, i. e. linear elastodynamics in 2D. As in our previous papers on linear elastostatics [13, 14], we adopt a true 2D setting where the third dimension is ignored [34], as opposed to e. g. a plane-stress or a plane-strain setting. This assumption is not critical at any point of the derivation of our LBM formulation, but just leads to simpler expressions; the extension to plane-stress or plane-strain conditions would be straightforward. Furthermore, we restrict ourselves to unbounded domains with periodicity or Dirichlet problems on rectangular domains. An extension towards Neumann boundary conditions and more general problem geometries will be the focus of future work. Next, we introduce the equivalent reformulation of our target problem as a first-order hyperbolic system of equations as an intermediate step to construct the vectorial LBM formulation. Lastly, all quantities are nondimensionalized for the handling in LBM.
@@ 2.1 Linear elastodynamics
$\bm{\sigma}:=\bar{\bm{\sigma}}/\rho=K/\rho(\nabla\cdot\bm{u})\bm{I}+\mu/\rho\left(\nabla\bm{u}+(\nabla\bm{u})^{T}-(\nabla\cdot\bm{u})\bm{I}\right)=:c_{K}^{2}(\nabla\cdot\bm{u})\bm{I}+c_{\mu}^{2}\left(\nabla\bm{u}+(\nabla\bm{u})^{T}-(\nabla\cdot\bm{u})\bm{I}\right).$ (3)
where we introduced the elastic wave speeds $c_{K}=\sqrt{K/\rho}$ and $c_{\mu}=\sqrt{\mu/\rho}$, involving the bulk and shear modulus $K,\mu>0$, respectively.
@@ 3.1 Vectorial lattice Boltzmann algorithm
$\mathbb{V}:=\{\bm{c}_{ij}\in\mathbb{R}^{2}:\bm{c}_{ij}=c(i\bm{e}_{x}+j\bm{e}_{y});\;\forall(i,j)\in\mathcal{V}\subset\mathbb{N}^{2}\},$ (21)
@@
The novel aspect of vectorial LBM is that the populations are vector-valued functions, i. e. $\bm{f}_{ij}:\mathbb{G}\times\mathbb{T}\rightarrow\mathbb{R}^{n}$ for all $(i,j)\in\mathcal{V}$ and with nonzero $n\in\mathbb{N}$ as the number of scalar components. Each of the scalar components of the population set is used to approximate the solution of one of the scalar equations in Eq. (13), which by itself is a simple hyperbolic conservation law-type problem. The number of microscopic velocities, i. e. $|\mathbb{V}|$, determines the amount of information and the complexity of the problem that can be solved by the LBM formulation because it governs the number of independent moments that can be computed from the available populations [35, 8, 23]. Because of the simplicity of the individual scalar problems, it turns out that a D2Q4 stencil involving four velocities is sufficient. Following the nomenclature in [25], we thus adopt a vectorial D2Q45 LBM stencil for our problem in Eq. (13) with $n=5$ scalar equations. The corresponding index set of the velocities and populations is therefore defined as follows
$\mathcal{V}:=\{(1,0),(0,1),(-1,0),(0,-1))\},$ (22)
@@
$\bm{f}_{ij}(\bm{x}+\bm{c}_{ij}\Delta t,t+\Delta t)=\omega\bm{f}_{ij}^{eq}\left(\bm{f}_{ij}(\bm{x},t);\tilde{\bm{B}}(\bm{x},t)\right)+(1-\omega)\bm{f}_{ij}(\bm{x},t)+\Delta t(2-\omega)W_{ij}\tilde{\bm{B}}(\bm{x},t)\qquad\forall(\bm{x},t)\in\mathbb{G}\times\mathbb{T},$ (23)
which is decomposed into the two following stages:
@@ 4.1 Second-order consistency in the domain interior
First, we investigate the second-order consistency of the formulation introduced in Section 3.2. During the proof using asymptotic analysis it will turn out that the second-order consistency in the domain interior critically depends on the accuracy of the initialization (see Section 3.3) and the boundary formulation (see Section 3.4.2). To this end, we define the notion of a second-order consistent initialization and a second-order consistent boundary formulation, which implies consistent behavior w. r. t. the physical boundary condition at leading order of the asymptotic expansion and zero initial boundary conditions at the next higher order. Accordingly, let us introduce all requirements in the following proposition:
@@ 5.2 Stability of vectorial LBM for elastodynamics with periodic boundaries
$||\bm{f}||^{2}_{\bm{P}}(t+\Delta t)=\sum_{\bm{y}\in\mathbb{G}}|\bm{P}\bm{f}(\bm{y},t+\Delta t)|^{2}=\sum_{\bm{y}\in\mathbb{G}}\sum_{(k,l)\in\mathcal{V}}|\bm{f}_{ij}(\bm{y},t+\Delta t)|^{2}_{\bm{k}_{ij}}=\sum_{\bm{x}\in\mathbb{G}}\sum_{(i,j)\in\mathcal{V}}|\bm{f}^{*}_{ij}(\bm{x},t)|^{2}_{\bm{k}_{ij}}=||\bm{f}^{*}||^{2}_{\bm{P}}(t),$ (74)
@@ 5.3 Stability of vectorial LBM for elastodynamics with Dirichlet boundary conditions
$\sum_{\bm{y}\in\mathbb{G}}\sum_{(k,l)\in\mathcal{V}\setminus\mathcal{D}_{\bm{y}}}|\bm{f}_{ij}(\bm{y},t+\Delta t)|^{2}_{\bm{k}_{kl}}=\sum_{\bm{x}\in\mathbb{G}}\sum_{(i,j)\in\mathcal{V}\setminus(-\mathcal{D}_{\bm{x}})}|\bm{f}^{*}_{ij}(\bm{x},t)|^{2}_{\bm{k}_{ij}}=\sum_{\bm{x}\in\mathbb{G}\setminus\mathbb{B}}\sum_{(i,j)\in\mathcal{V}}|\bm{f}^{*}_{ij}(\bm{x},t)|^{2}_{\bm{k}_{ij}}+\sum_{\bm{x}\in\mathbb{B}}\sum_{(i,j)\in\mathcal{V}\setminus(-\mathcal{D}_{\bm{x}})}|\bm{f}^{*}_{ij}(\bm{x},t)|^{2}_{\bm{k}_{ij}}.$ (78)
@@
$\begin{split}||\bm{f}||^{2}_{\bm{P}}(t+\Delta t)&=\sum_{\bm{y}\in\mathbb{G}}\sum_{(i,j)\in\mathcal{V}\setminus\mathcal{D}_{\bm{y}}}|\bm{f}_{ij}(\bm{y},t+\Delta t)|^{2}_{\bm{k}_{ij}}+\sum_{\bm{y}\in\mathbb{B}}\sum_{(i,j)\in\mathcal{D}_{\bm{x}}}|\bm{f}_{ij}(\bm{x},t+\Delta t)|^{2}_{\bm{k}_{ij}}\\ &=\sum_{\bm{x}\in\mathbb{G}\setminus\mathbb{B}}\sum_{(i,j)\in\mathcal{V}}|\bm{f}^{*}_{ij}(\bm{x},t)|^{2}_{\bm{k}_{ij}}+\sum_{\bm{x}\in\mathbb{B}}\sum_{(i,j)\in\mathcal{V}\setminus(-\mathcal{D}_{\bm{x}})}|\bm{f}^{*}_{ij}(\bm{x},t)|^{2}_{\bm{k}_{ij}}+\sum_{\bm{x}\in\mathbb{B}}\sum_{(\bar{\imath},\bar{\jmath})\in(-\mathcal{D}_{\bm{x}})}|\bm{f}^{*}_{\bar{\imath}\bar{\jmath}}(\bm{x},t)|^{2}_{\bm{k}_{\bar{\imath}\bar{\jmath}}}=||\bm{f}^{*}||^{2}_{\bm{P}}(t),\end{split}$ (81)
@@ 6.1 Convergence studies
$\hat{\bm{u}}=\left[\begin{matrix}\sin{4\pi(x-0.3t)}\cos{2\pi(y-0.8t)}\sin{4\pi(t-0.1)}\\ \cos{4\pi(x-0.7t)}\sin{2\pi(y-0.1t)}\cos{4\pi(t+0.4)}\end{matrix}\right]\qquad\forall(\bm{x},t)\in\bar{\Omega}\times[0,\tilde{t}_{f}),$ (90)
@@
For better comparability, the perform the exact same analysis, but only replace the periodicity by Dirichlet boundary conditions on the whole boundary. The results of the convergence studies are summarized in Figures 6 and 7. Once again, we obtain second-order convergence of the error in the displacement solution in both norms, which is a strong confirmation of the analysis. Further, we can notice a slightly higher $\text{Linf}\left(\bm{e}^{\bm{u}}\right)$ when compared to the periodic case, which can be attributed to additional error contributions due to the boundary formulation. The error of the Cauchy stress solution given in Figure 7 now shows a different behavior. Whereas the L2-norm still shows approximately second-order convergence, we observe a reduced (first-order) convergence in the Linf-norm. This lets us conclude that the Dirichlet boundary formulation introduces a first-order error in the Cauchy stress, which remains localized at the boundary. Note that for the material parameters $(\tilde{c}_{K}^{2},\tilde{c}_{\mu}^{2})\in\{(1.5,0.0),(1.4,0.1)\}$ the convergence order degrades only for relatively fine discretizations to first order – probably because other second-order errors with a higher constant overshadow the first-order error for finite $\Delta x$ (or equivalently $\Delta t$). A similar behavior was already observed with Dirichlet boundary conditions in the linear elastostatics case and was extensively analyzed in [14]. Overall, also with Dirichlet boundary conditions, we retain the second-order convergence in both solution fields in the L2 norm, i. e. in an average sense. Note that, for similar problems, the finite element method with linear ansatz functions achieves second-order convergence for the displacement solution but only first-order convergence for the Cauchy stress solution [48].
@@ 6.2 Numerical stability tests
$\bm{u}_{0}(\bm{x})=\left[\begin{matrix}\sin{4\pi x}\sin{2\pi y}\sin{4\pi(t-0.1)}\\ \sin{4\pi x}\sin{2\pi y}\sin{4\pi(t+0.4)}\end{matrix}\right],\quad\bm{v}_{0}(\bm{x})=\partial_{t}\bm{u}_{0}(\bm{x})\qquad\forall\bm{x}\in\Omega,$ (91)
@@ 7 Conclusions
With this initial contribution, we obtained a novel LBM formulation for linear elastodynamics that overcomes the shortcomings of previous works – by providing second-order consistency and rigorous stability estimates. The main limitation of the present formulation is that it can only be used for periodic or Dirichlet problems with grid-aligned boundaries. Therefore, the next natural step will be to generalize the boundary formulation in terms of both Neumann boundary conditions and arbitrary geometries. We also believe that the new method deserves a thorough comparison with existing competing methods, and that it holds a great potential for further developments in several directions, which will be the focus of future work.

### v2

@@ 1 Introduction
To circumvent the aforementioned issues, in this paper we adopt the so-called vectorial LBM, which involves vector-valued instead of scalar populations [24, 25, 26, 27, 28, 29]. In a sense, multi distribution formulations such as those in [30, 31] can be also subsumed into the class of vectorial LBM. If we consider the so-called acoustic scaling, it is straightforward to show that the vectorial LBM with a vectorial counterpart of the well-known BGK collision operator approximates first-order hyperbolic systems of PDEs with second-order consistency in the non-dissipative limit [25]. In fact, the vectorial LBM can be seen as a particularly simple numerical approximation of so-called relaxation schemes for systems of conservation laws [32, 33].
@@ 2 Linear elastodynamics and its reformulation for LBM
In this section, we start by introducing our target problem, i. e. linear elastodynamics in 2D. We will show that it is possible to handle both a pure 2D setting [34] and 2D domains embedded in 3D space with the plane-strain or plane-stress assumption by modifying the definition of a single problem parameter. As a result, the entire analysis can be applied to all cases without any further changes. In this contribution, we restrict ourselves to unbounded domains with periodicity or Dirichlet problems on rectangular domains; an extension towards Neumann boundary conditions and more general problem geometries will be the focus of future work. Next, we introduce the equivalent reformulation of our target problem as a first-order hyperbolic system of equations as an intermediate step to construct the vectorial LBM formulation. Lastly, all quantities are nondimensionalized for the handling in LBM.
@@ 2.1 Linear elastodynamics
$\bm{\sigma}:=\bar{\bm{\sigma}}/\rho=c_{K}^{2}(\nabla\cdot\bm{u})\bm{I}+c_{\mu}^{2}\left(\nabla\bm{u}+(\nabla\bm{u})^{T}-(\nabla\cdot\bm{u})\bm{I}\right).$ (3)
Here we introduced the elastic wave speeds $c_{K}$ and $c_{\mu}$ given by the following expressions
2D continua: $c_{K}:=\sqrt{K/\rho}$ (4)
Plane strain: $c_{K}:=\sqrt{\frac{3K+\mu}{3\rho}}$ (5)
Plane stress: $c_{K}:=\sqrt{\frac{9K\mu}{(3K+4\mu)\rho}}$ (6)
All cases: $c_{\mu}:=\sqrt{\mu/\rho},$ (7)
with $K,\mu>0$ as the bulk and shear modulus, respectively.
@@ 3.1 Vectorial lattice Boltzmann algorithm
$\mathbb{V}:=\{\bm{c}_{ij}\in\mathbb{R}^{2}:\bm{c}_{ij}=c(i\bm{e}_{x}+j\bm{e}_{y});\;\forall(i,j)\in\mathcal{V}\subset\mathbb{Z}^{2}\},$ (25)
@@
The novel aspect of vectorial LBM is that the populations are vector-valued functions, i. e. $\bm{f}_{ij}:\mathbb{G}\times\mathbb{T}\rightarrow\mathbb{R}^{n}$ for all $(i,j)\in\mathcal{V}$ and with nonzero $n\in\mathbb{N}$ as the number of scalar components. Each of the scalar components of the population set is used to approximate the solution of one of the scalar equations in Eq. (17), which by itself is a simple hyperbolic conservation law-type problem. The number of microscopic velocities, i. e. $|\mathbb{V}|$, determines the amount of information and the complexity of the problem that can be solved by the LBM formulation because it governs the number of independent moments that can be computed from the available populations [35, 8, 23]. In our case, each scalar equation in Eq. (17) involves three quantities (one component of $\bm{U}$, $\bm{\Phi}_{x}$ and $\bm{\Phi}_{y}$), therefore we need at least three independent moments and velocities per scalar equation. In order to obtain a regular Cartesian lattice in 2D, we add one additional velocity and set the corresponding extra moment to zero (see Eq. (30) in Section 3.2). Following the nomenclature in [25], we thus adopt a vectorial D2Q45 LBM stencil for our problem in Eq. (17) with $n=5$ scalar equations. The corresponding index set of the velocities and populations is therefore defined as follows
$\mathcal{V}:=\{(1,0),(0,1),(-1,0),(0,-1)\},$ (26)
@@
$\bm{f}_{ij}(\bm{x}+\bm{c}_{ij}\Delta t,t+\Delta t)=\omega\bm{f}_{ij}^{eq}\left(\bm{f}_{ij}(\bm{x},t);\tilde{\bm{B}}(\bm{x},t)\right)+(1-\omega)\bm{f}_{ij}(\bm{x},t)+\Delta t(2-\omega)W_{ij}\tilde{\bm{B}}(\bm{x},t)\qquad\forall(\bm{x},t)\in\mathbb{G}\times\mathbb{T}.$ (27)
The origin of this equation from gas dynamics and the main steps needed to obtain this numerical algorithm are summarized in A. Conceptually, Eq. (27) is decomposed into the two following stages:
@@ 4.1 Second-order consistency in the domain interior
First, we investigate the second-order consistency of the formulation introduced in Section 3.2. During the proof using asymptotic analysis it will turn out that the second-order consistency in the domain interior critically depends on the accuracy of the initialization (see Section 3.3) and the boundary formulation (see Section 3.4.2). To this end, we define the notion of a second-order consistent initialization and a second-order consistent boundary formulation, which implies consistent behavior with respect to the physical initial or boundary condition at leading order of the asymptotic expansion and zero initial or boundary conditions at the next higher order. Accordingly, let us introduce all requirements in the following proposition:
@@ 5.2 Stability of vectorial LBM for elastodynamics with periodic boundaries
$||\bm{f}||^{2}_{\bm{P}}(t+\Delta t)=\sum_{\bm{y}\in\mathbb{G}}|\bm{P}\bm{f}(\bm{y},t+\Delta t)|^{2}=\sum_{\bm{y}\in\mathbb{G}}\sum_{(i,j)\in\mathcal{V}}|\bm{f}_{ij}(\bm{y},t+\Delta t)|^{2}_{\bm{k}_{ij}}=\sum_{\bm{x}\in\mathbb{G}}\sum_{(i,j)\in\mathcal{V}}|\bm{f}^{*}_{ij}(\bm{x},t)|^{2}_{\bm{k}_{ij}}=||\bm{f}^{*}||^{2}_{\bm{P}}(t),$ (78)
@@ 5.3 Stability of vectorial LBM for elastodynamics with Dirichlet boundary conditions
$\sum_{\bm{y}\in\mathbb{G}}\sum_{(i,j)\in\mathcal{V}\setminus\mathcal{D}_{\bm{y}}}|\bm{f}_{ij}(\bm{y},t+\Delta t)|^{2}_{\bm{k}_{ij}}=\sum_{\bm{x}\in\mathbb{G}}\sum_{(i,j)\in\mathcal{V}\setminus(-\mathcal{D}_{\bm{x}})}|\bm{f}^{*}_{ij}(\bm{x},t)|^{2}_{\bm{k}_{ij}}=\sum_{\bm{x}\in\mathbb{G}\setminus\mathbb{B}}\sum_{(i,j)\in\mathcal{V}}|\bm{f}^{*}_{ij}(\bm{x},t)|^{2}_{\bm{k}_{ij}}+\sum_{\bm{x}\in\mathbb{B}}\sum_{(i,j)\in\mathcal{V}\setminus(-\mathcal{D}_{\bm{x}})}|\bm{f}^{*}_{ij}(\bm{x},t)|^{2}_{\bm{k}_{ij}}.$ (82)
@@
$\begin{split}||\bm{f}||^{2}_{\bm{P}}(t+\Delta t)&=\sum_{\bm{y}\in\mathbb{G}}\sum_{(i,j)\in\mathcal{V}\setminus\mathcal{D}_{\bm{y}}}|\bm{f}_{ij}(\bm{y},t+\Delta t)|^{2}_{\bm{k}_{ij}}+\sum_{\bm{x}\in\mathbb{B}}\sum_{(i,j)\in\mathcal{D}_{\bm{x}}}|\bm{f}_{ij}(\bm{x},t+\Delta t)|^{2}_{\bm{k}_{ij}}\\ &=\sum_{\bm{x}\in\mathbb{G}\setminus\mathbb{B}}\sum_{(i,j)\in\mathcal{V}}|\bm{f}^{*}_{ij}(\bm{x},t)|^{2}_{\bm{k}_{ij}}+\sum_{\bm{x}\in\mathbb{B}}\sum_{(i,j)\in\mathcal{V}\setminus(-\mathcal{D}_{\bm{x}})}|\bm{f}^{*}_{ij}(\bm{x},t)|^{2}_{\bm{k}_{ij}}+\sum_{\bm{x}\in\mathbb{B}}\sum_{(\bar{\imath},\bar{\jmath})\in(-\mathcal{D}_{\bm{x}})}|\bm{f}^{*}_{\bar{\imath}\bar{\jmath}}(\bm{x},t)|^{2}_{\bm{k}_{\bar{\imath}\bar{\jmath}}}=||\bm{f}^{*}||^{2}_{\bm{P}}(t),\end{split}$ (85)
@@ 6.1 Convergence studies
$\hat{\bm{u}}=\left[\begin{matrix}\sin{(4\pi(x-0.3t))}\cos{(2\pi(y-0.8t))}\sin{(4\pi(t-0.1))}\\ \cos{(4\pi(x-0.7t))}\sin{(2\pi(y-0.1t))}\cos{(4\pi(t+0.4))}\end{matrix}\right]\qquad\forall(\bm{x},t)\in\bar{\Omega}\times[0,\tilde{t}_{f}),$ (94)
@@
For better comparability, we perform the exact same analysis, but only replace the periodicity by Dirichlet boundary conditions on the whole boundary. The results of the convergence studies are summarized in Figures 6 and 7. Once again, we obtain second-order convergence of the error in the displacement solution in both norms, which is a strong confirmation of the analysis. Further, we can notice a slightly higher $\text{Linf}\left(\bm{e}^{\bm{u}}\right)$ when compared to the periodic case, which can be attributed to additional error contributions due to the boundary formulation. The error of the Cauchy stress solution given in Figure 7 now shows a different behavior. Whereas the L2-norm still shows approximately second-order convergence, we observe a reduced (first-order) convergence in the Linf-norm. This lets us conclude that the Dirichlet boundary formulation introduces a first-order error in the Cauchy stress, which remains localized at the boundary. Note that for the material parameters $(\tilde{c}_{K}^{2},\tilde{c}_{\mu}^{2})\in\{(1.5,0.0),(1.4,0.1)\}$ the convergence order degrades only for relatively fine discretizations to first order – probably because other second-order errors with a higher constant overshadow the first-order error for finite $\Delta x$ (or equivalently $\Delta t$). A similar behavior was already observed with Dirichlet boundary conditions in the linear elastostatics case and was extensively analyzed in [14]. Overall, also with Dirichlet boundary conditions, we retain the second-order convergence in both solution fields in the L2 norm, i. e. in an average sense. Note that, for similar problems, the finite element method with linear ansatz functions achieves second-order convergence for the displacement solution but only first-order convergence for the Cauchy stress solution [48].
@@ 6.2 Numerical stability tests
$\bm{u}_{0}(\bm{x})=\left[\begin{matrix}\sin{(4\pi x)}\sin{(2\pi y)}\sin{(4\pi(t-0.1))}\\ \sin{(4\pi x)}\sin{(2\pi y)}\sin{(4\pi(t+0.4))}\end{matrix}\right],\quad\bm{v}_{0}(\bm{x})=\partial_{t}\bm{u}_{0}(\bm{x})\qquad\forall\bm{x}\in\Omega,$ (95)
@@ 7 Conclusions
With this initial contribution, we obtained a novel LBM formulation for linear elastodynamics that overcomes the shortcomings of previous works – by providing second-order consistency and rigorous stability estimates. Furthermore, note that the definition of $\bm{\Phi}_{x}$ and $\bm{\Phi}_{y}$ in Eq. (17) can be generalized to arbitrary nonlinear behavior as well and therefore a straightforward extension to nonlinear material behavior and/or large deformation settings appears feasible. In this contribution, linearity of the problem has only been exploited in the consistency analysis and construction of the stability proof. In consistency analysis, nonlinear behavior can be easily accommodated, whereas the stability analysis potentially requires a more general treatment for nonlinear problems.
The main limitation of the present formulation is that it can only be used for periodic or Dirichlet problems with grid-aligned boundaries. Therefore, the next natural step will be to generalize the boundary formulation in terms of both Neumann boundary conditions and arbitrary geometries. Already in the context of ”conventional” LBM for fluid mechanics or elastostatics [13, 14], the design of second-order accurate boundary formulations on arbitrary domains turns out to be challenging. The main reasons include
  • 1.
    mismatch of the number of physical boundary conditions present and the number of missing incoming populations that need to be reconstructed [49, 50];
  • 2.
    the component-wise resolution of vector or tensor-valued quantities in LBM making rotation into a boundary-aligned reference system challenging [51, 14];
  • 3.
    compensation of higher-order derivatives that naturally appear during the asymptotic expansion of boundary formulations [52, 14]
  • 4.
    the intention of preserving a local formulation that requires only minimal neighbor node information [38, 39];
  • 5.
    stability issues that only in few specialized cases have been analyzed analytically (such as in this work for aligned boundaries and in [38, 43, 45]).
Accordingly, the derivation of boundary formulations on arbitrary geometries as well as Neumann-type boundary conditions will require extensive additional investigations and will be the focus of future studies.
Another limitation of the method is that by coupling the velocity discretization to the spatial discretization the LBM is restricted to Cartesian grids. While this is a limitation of the LBM irrespective of the equation being solved, there are several methods known from fluid dynamics to overcome it. Among these are interpolation based methods [53] and semi-Lagrangian off-lattice methods [54] that modify the streaming step. Besides these there exist many established grid refinement methods that couple Cartesian grids with varying resolution [55, 56, 57, 58, 7]. All these method use interpolation of populations or moments at the interface together with an appropriate reconstruction of the non-equilibrium populations. The so-called compact interpolation methods [58, 7] acquire the interpolation polynomial by including the gradient information from asymptotic analysis. A similar approach would also be applicable for our vectorial LBM, which however requires further detailed investigations.